40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 27 04:11:21 2015, Jenne, Update, LSC, Small tweaks to the locking CARM_27Jan2015_JCD.pdfDARM_27Jan2015.pdf
    Reply  Tue Jan 27 15:45:26 2015, diego, Update, LSC, Small tweaks to the locking 
Message ID: 10942     Entry time: Tue Jan 27 04:11:21 2015     Reply to this: 10944
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Small tweaks to the locking 

[Jenne, Diego, EricQ]

We did a series of small things that may have helped with the locking, although we didn't actually get anywhere closer in CARM offset.

  • Removed the demodulation phase from the UGF servos.
    • We don't care about the phase value, just the magnitude.
    • Since we are using these through transitions between error signals, as well as with different filters coming on and off, the demodulation phase isn't constant, so the UGF servo is getting the wrong answer, and was throwing us out of lock.
    • The problems that Rana and I saw last time we had the sum of the sqrt were later discovered to be attributable to losing the peak in the noise, and hitting saturation limiters in the model, so not the fault of the sum of the squares.
    • I don't actually take the square root.  As Koji pointed out, the very next thing that happens to the signal is mag2dB, which is usually 20*log10(mag).  To compensate for the fact that I'm not taking the square root, I just use 10*log10(mag).  This removes one element of non-linearity, and leaves it at about the same number of square-ings as the complex division.
    • The excitation and measurement still happen after the multiplication.
    • The UGF screens will be updated tomorrow to reflect the change.
  • Added the new error signal rows to the LSC model's DAQ list.  So, now DARM_A_ERR and DARM_B_ERR are both acquired (and the same for CARM, MICH and PRCL).  This is to allow us to look at _DQ channels with dataviewer and DTT without having to clear the testpoints all the time.
    • We were running into too many channels being recorded, so we are not keeping the SRCL A & B signals right now.  Also, the CESAR signals that are no longer in use have been removed from the DAQ list.
  • Added a comb to the ASDC and POPDC signals, to remove the 60Hz harmonics from the MICH error signal.
    • The harmonics of the 60Hz line were dominating by more than an order of magnitude the RMS for the MICH control signal.  We couldn't afford too much phase at a few tens of Hz, so we do not notch out the original 60Hz line, although it isn't as big as, say, the 180Hz line.  So, I think that the notches are for the 2nd, 3rd, 4th and 5th order harmonics of 60Hz.  This significantly improved the RMS of the MICH output signal.
  • Lowered the MICH UGF slightly, from 48Hz to 41Hz. 
    • We wanted to go lower, to maybe 30Hz-ish, but we don't have the phase margin.  The roll mode notch is in the way, so we compromised at 41Hz.  With the comb mentioned above, the MICH control signal looks much more reasonable, and we're not injecting oodles of sensing noise into the BS.
    • Together with the comb, this ensures that we are not constantly railing the MICH output limiter, which lost us lock several times.
  • Increased the POPDC analog whitening gain from 0dB to 18dB.  We will certainly saturate POPDC when the carrier is resonant, but we were hoping to improve our SNR in our MICH error signal.  It helped noticeably, although we'll have to think about the fact that if it saturates while we're trying to acquire, the AS/POP composite signal won't be any good, and we'll kick the optics.  Also, we may have to lower the gain again as the carrier comes in to resonance.  Anyhow, something to think about.  Right now the gain is left at 18dB, and the dark offsets are set to match.
  • We may have been using the wrong side of the MICH offset, according to Q's plots.  We determined tonight that we want a (-) in the MICH gain, although the offset values can stay the same.  Even though we tried this, we didn't really get any farther in CARM offset reduction.
  • We took 2 sets of CARM and DARM loops.  They are both at 50% MICH offset, although I don't remember which sign the first one had.  The second one, at arm powers of 1.4, definitely had the new negative MICH gain. 
    • The first loop was taken at 50% MICH offset (don't remember which sign), and arm powers of about 1.15.
    • The second loop was taken at 50% MICH offset with negative gain, and arm powers of about 1.4.
    • While these numbers are not so different, maybe the first one was at roughly 300pm, and the second was at roughly 200pm, the loop shapes change dramatically.
    • The phase goes flat and we lose some of the phase margin.  Also, the magnitude is starting to get wiggly at lower frequencies.
    • See the transfer functions attached below.
  • While we were sitting at about arm powers of 1.4, with the 50% MICH offset with the negative sign in the gain, we set the REFL 11 demod phase.  It was 18.9deg (I don't remember from when), but we set it to 2.9deg to minimize the PRCL actuation in the Q-phase.  Oscillator was at 443Hz, about 10 counts. 
    • The coherence between the sqrtInvTrans signal and REFL11I looked pretty good from a few Hz to a few tens of Hz. 
    • It was late, so we decided to try transitioning over to REFL11I, and failed at the first very baby step.  So.  Ooops.
    • We should look more carefully at the TF between our current CARM signal and REFL11I.
    • We should also seriously consider using a normalized RF signal.  The SNR in the transmission PDs is just fine, although POPDC isn't a perfect choice at such high MICH offsets.
Attachment 1: CARM_27Jan2015_JCD.pdf  13 kB  | Hide | Hide all
CARM_27Jan2015_JCD.pdf
Attachment 2: DARM_27Jan2015.pdf  9 kB  | Hide | Hide all
DARM_27Jan2015.pdf
ELOG V3.1.3-